OpenStudy (anonymous):

Solve using factoring. 2x^2 +6x = 0 ^ = raised to, ex. 2 squared

OpenStudy (anonymous):

x(2x+6) = 0 x=0 or 2x+6 = 0 2x =-6 x = -6/2 x = -3

OpenStudy (anonymous):

that would help. but im trying to use factoring to find two points

OpenStudy (anonymous):

like (5,2) and (3,5)

OpenStudy (anonymous):

the roots?

OpenStudy (anonymous):

uh, yeah

OpenStudy (anonymous):

then in the form (x,0) it would be (0,0) and (-3,0)

OpenStudy (anonymous):

that works, thank you

OpenStudy (anonymous):

np